OpenStudy (anonymous):

A video game has been developed whereby the player must earn a certain number of experience points before progressing to the next level. On level 1, a player must earn 1 experience point. On level 2, a player must earn 4 experience points. On level 3, a player must earn 8 experience points. On level 4, a player must earn 13 experience points. Which recursive equation represents the pattern?

OpenStudy (anonymous):

a. an = ((an - 1)) + n b. an = ((an - 1)) + (n + 1) c. an = ((an - 1)) + 2n d. an = ((an - 1)) + (2n + 1)

OpenStudy (anonymous):

@Skyz halp pls

OpenStudy (skyz):

o.o flutter dat i aint got time for dat xD

OpenStudy (anonymous):

D; halp pls crys

OpenStudy (skyz):

i think c... honestly because of how i plugged in the shiz

OpenStudy (skyz):

but yeah idek this

OpenStudy (anonymous):

well educated guess ftw